Reports 1-1 of 1 Clear search Modify search
fabian.arellano - 17:32 Thursday 08 April 2021 (16427) Print this report
Comment to Tightening of the inner wire locating clamps. (16417)

With Hirata-san.

Summary: we finished the assembly of the RM and the IRM and released the payload and BF. There seems to be a large amount of positive pitch in the IM according to the position of the OSEM flags. We will aim to correct it with picomotors.

See pictures in album PR2 Remedying Work.

  • We put back onto the RM the coil bodies. We checked that the efective length of the push screws were 8 mm.
  • We removed the spacers (tips of PEEK cable ties) that were in between the RM sides the the clamped wires, just underneath the wire breakers.
  • We released the mirror and the RM to let the RM hang free.
  • We inspected the outer wire clamps that we had removed. See pictures in the album. The scars were more superficial than in PR3 becuase the clamps had not been tightened properly before.
  • We put the outer clamp in place, made sure the tungsten wires were witin the grooves and tightened the screws with a 2 Nm torque.
  • At  the mirror, we took pictures of the magnet standoffs position within their respective cavities. It seems they are roughy at the same place as before but a more detailed comparison would be good.
  • We put magnets on the H2 and H3 OSEM pedestals. North pole out.
  • We put back the IRM side panels.
  • We put the OSEM flags in place and connected the OSEMs.
  • Released the mirror, RM, IM, IRM and BF.
  • According to the position of the OSEM flags there seems to be a lot of positive pitch. We will aim to correct it with the picomotors.
Search Help